# GraphBEV: Towards Robust BEV Feature Alignment for Multi-Modal 3D Object Detection

### v2

@@ 1 Introduction
Figure 1: (a) Feature misalignment primarily arises from overlooking the projection matrix errors between LiDAR and camera, leading to LiDAR-to-camera providing inaccurate depth for surrounding neighbors. (b) We propose GraphBEV, enhancing LiDAR-to-camera projected depth with neighboring depths through graph-based neighbor information construction, for enriched contextual depth feature learning. Subsequently, we achieve the alignment of global multi-modal features by simulating the offsets of LiDAR and camera BEV features, and subsequently employing learnable offsets. (c) Empirical results reveal that our GraphBEV surpasses the BEVFusion [34] on the nuScenes by a margin of 1.6% mAP on nuScenes validation dataset [4] and by over 8.3% on noisy misalignment settings [12].
@@
3D object detection is a critical component of autonomous driving systems, designed to accurately identify and localize objects such as cars, pedestrians, and other elements within a 3D environment [49, 58]. For robust and high-quality detection, current practices predominantly adhere to a multi-modal fusion paradigm like BEVFusion[34, 29]. Different modalities often provide complementary information. For instance, images contain rich semantic representations, yet lack depth information. In contrast, point clouds offer geometric and depth information but are sparse and devoid of semantic information. Therefore, effectively leveraging the advantages of multi-modal data while mitigating their limitations is crucial for enhancing the robustness and accuracy of perception systems [58].
@@
Feature misalignment is a significant challenge in practical applications of multi-modal 3D object detection, primarily caused by calibration matrix errors between LiDAR and camera sensors [70, 12, 49], as shown in Figure 1(a). Multi-modal 3D object detection has evolved from early point-level [47, 55, 17, 33, 56, 68] and feature-level [52, 10, 2, 9, 50, 48] methods to the currently prevalent methods of Bird’s Eye View (BEV) fusion like BEVFusion [29, 34]. Although effective on clean datasets like nuScenes [4], the performance of BEVFusion[34] deteriorates on misaligned data, as illustrated in Figure 1(c). This performance decline is primarily due to calibration errors between LiDAR and camera, exacerbated by factors like road vibrations [70]. These inherent errors cannot be corrected through online calibration, presenting a significant challenge[70, 12].
@@
Most feature-level multi-modal methods [24, 10, 1, 66, 72] employ Cross Attention to query features of a specific modality, circumventing the use of projection matrices. A few feature-level multi-modal methods [69, 50, 48, 9, 23, 71, 22, 19] have sought to mitigate these errors through projection offsets or neighboring projections. A few BEV-based methods, such as ObjectFusion [5], completely get rid of camera-to-BEV transformation during fusion to align object-centric features across different modalities. MetaBEV [14] utilizes Cross Deformable Attention for feature misalignment but overlooks depth estimation errors in view transformation, aligning features only during LiDAR and camera BEV fusion.
@@
BEVFusion[34, 29] unites camera and LiDAR data in BEV space, enhancing detection but overlooks feature misalignment in real-world applications. It is primarily evident in two aspects: 1) BEVFusion[34] transforms multi-image features into a unified BEV representation using BEVDepth’s[26] explicit depth supervision from LiDAR-to-camera. While this LiDAR-to-camera strategy offers more reliable depth than LSS[40], it overlooks the misalignment between LiDAR and camera in real-world scenarios, leading to local misalignment. 2) In the LiDAR-camera BEV fusion, misalignment of BEV features due to depth inaccuracies is overlooked by directly concatenating representations and applying basic convolution, as described in the text of BEVFusion[34], leading to global misalignment.
@@
In this research, we propose a robust fusion framework, named GraphBEV, to address the above feature misalignment problems. To tackle the local misalignment issue in camera-to-BEV transformation, we propose a LocalAlign module that first obtains neighbor depth information through a Graph in the View Transformation step of BEVFusion’s [34] camera Branch, based on the explicit depth provided by the LiDAR-to-camera projection. Subsequently, we propose a GlobalAlign module that encodes the projected depth from LiDAR-to-camera and the neighboring depth through dual depth encoding to generate a new Reliable depth representation incorporating neighbor information. Furthermore, to address the global misalignment problem in the fusion of LiDAR and camera BEV features, we effectively resolve the global misalignment between the two BEV feature representations by dynamically generating offsets. To validate the effectiveness of GraphBEV, we evaluated it on the nuScenes dataset, a well-known benchmark for 3D object detection. Our GraphBEV not only achieves state-of-the-art (SOTA) performance under clean settings, but also outperforms BEVFusion by over 8.3% on non-aligned noisy settings nuScenes dataset provided by Dong et al [12]. It is noteworthy that our GraphBEV experiences only a slight performance drop under noisy settings compared to clean settings. Notably, our GraphBEV bolsters the real-world applicability of BEV-based methods [34, 29] by addressing sensor calibration errors. Our contributions are summarized as below:
@@
    By deeply analyzing the fundamental causes of feature misalignment, we propose LocalAlign and GlobalAlign modules within our GraphBEV to address local misalignments from imprecise depth and global misalignments between LiDAR and camera BEV features.
@@
    Extensive experiments validate the effectiveness of our GraphBEV, demonstrating competitive performance on nuScenes. Notably, GraphBEV maintains comparable performance across both clean settings and misaligned noisy conditions.
@@ 2.3 Multi-modal 3D Object Detection
Multi-modal 3D object detection refers to using data features from different sensors and integrating these features to achieve complementarity, thus enabling the detection of 3D objects. Previous multi-model methods can be coarsely classified into three types by fusion, i.e., point-level, feature-level, and BEV-based methods. Point-level [47, 55, 17, 33, 56, 68] and feature-level [52, 10, 2, 9, 50, 48] typically leverage image features to augment LiDAR points or 3D object proposals. BEV-based methods [34, 29, 14, 5] efficiently unify the representations of LiDAR and camera into BEV space. Although BEVFusion[34, 29] achieves high performance, they are typically tested on clean datasets like nuScenes[4], overlooking real-world complexities, especially feature misalignment, which hampers their application.
@@ 3 Method
Figure 2: An overview of GraphBEV framework. The LiDAR branch almost follows the baselines  [34, 1] to generate LiDAR BEV features. In the camera branch, first, we extract camera BEV features using proposed LocalAlign module that aim to addressing local misalignment due sensor calibration errors. Subsequently, we simulate the offset noisy of LiDAR and camera BEV features, followed by aligning global multi-modal features through learnable offsets. It is noteworthy that we only add offset noise to the GlobalAlign module during training to simulate global misalignment issues. Finally, we employ a dense detection head [1] to accomplish the 3D detection task.
To address the feature misalignment issue in previous BEV-based methods [34, 29], we propose a robust fusion framework, named GraphBEV. We provide an overview of our framework in Figure 2. Taking input from different sensors, including LiDAR and camera, we first apply modality-specific encoders, Swin-Transformer [32] as camera encoder and Second[61] as LiDAR encoder, to extract their respective features. Then, the camera features are transformed into the camera BEV feature through our proposed LocalAlign module, aiming to mitigate the local misalignment caused by the projection error between LiDAR and the camera in the camera-to-BEV process of previous BEV-based methods [34, 29]. The LiDAR features are then compressed along the Z-axis to represent 3D features as 2D LiDAR BEV features. Next, we propose a GlobalAlign module that further alleviates global misalignment between different modalities, including LiDAR and camera BEV features. Finally, we append detection head [1] to accomplish the 3D detection task. Our baseline is the BEVFusion[34], within which we detail the introduction of the LocalAlign and GlobalAlign modules below.
@@ 3.1 LocalAlign
Figure 3: An overview of LocalAlign pipline. The LocalAlign module addresses Local Misalignment from LiDAR-to-camera by enhancing camera-to-BEV Transform with neighboring depth features using a KD-Tree algorithm for nearest-neighbor relations.
To facilitate the transformation of camera features into BEV features, BEVFusion [34, 29] adopts the LSS-based[40] methods like BEVDepth [26] leverages LiDAR-to-camera to provide projected depth, thereby enabling the fusion of depth and image features. In the process of camera-to-BEV, BEVFusion [34] and BEVDepth [26] operate under the assumption that the depth information provided by LiDAR-to-camera projection is accurate and reliable. However, they overlooks the complexities inherent in real-world scenarios, where most of the projection matrices between LiDAR and cameras are calibrated manually. Such calibration inevitably introduces projection errors, leading to depth misalignment—where the depths of surrounding neighbors are projected as the pixel’s depth. This depth misalignment results in inaccuracies within the depth features, causing misalignment during the transformation of multi-view into BEV representations. Given that LSS-based methods [40, 26] rely on depth estimation from pixel-level features with inaccuracies in detailing, this leads to local misalignment within camera BEV features. This underscores the challenges of ensuring precise depth estimation within the BEVFusion [34] and highlights the importance of robust methods to address projection errors.
@@
Input: The indices of the projected pixels $M_{\text{Coords}}\in\mathbb{R}^{N_{P}\times 2}$. Hyper-parameter: Number of neighbors $K_{\text{graph}}=8$.
@@ 2       Function KD-Tree ($M_{\text{Coords}}$, ${M_{\text{Coords}}}_{i}$, $K_{\text{graph}}$)
3             Compute the Euclidean distance between ${M_{\text{Coords}}}_{i}$ and $M_{\text{Coords}}$ Indices = argsort(distances) return $M_{\text{Coords}}[1:K_{\text{graph}}]$
@@
6             Neighbors = KD-Tree($M_{\text{Coords}}$, ${M_{\text{Coords}}}_{i}$, $K_{\text{graph}}$) ${M_{\text{Coords}}}_{i}$ = Neighbors
@@
We employ the KD-Tree algorithm to obtain the indices of the projected pixels’ neighbors, defined as $M_{K_{\text{Coords}}}\in\mathbb{R}^{N_{P}\times K_{\text{graph}}\times 2}$, where $K_{\text{graph}}$ denotes the number of neighbors for each projected pixel. The process is outlined in Algorithm (1). It is worth noting that we simplified the process of the KD-Tree algorithm, and the code can be referred to in scipy [^1] https://github.com/minrk/scipy-1/blob/master/scipy/spatial/ckdtree.c. Then, we obtain the surrounding neighbor depth $D_{K}\in\mathbb{R}^{B_{S}\times N_{C}\times K_{\text{graph}}\times H\times W}$ by indexing $D_{S}$ with $M_{\text{Coords}}$. Then, $D_{S}$ and $D_{K}$ simultaneously enter the Dual Transform module for deep feature encoding. The shapes of $D_{S}$ and $D_{K}$ are respectively modified to $[B_{S}\times N_{\text{Cam}},1,H,W]$ and $[B_{S}\times N_{\text{Cam}},K_{\text{graph}},H,W]$ before being fed into the Dual Transform module. This module comprises straightforward components, including convolutional layers, Batch Normalization, and ReLU activations, as illustrated in Figure 3. The outcome of this process is the Dual Depth feature, denoted as $D_{\text{SK}}$, and its shape is $[B_{S}\times N_{\text{Cam}},C_{\text{SK}},\frac{H}{8},\frac{W}{8}]$. The camera Encoder outputs multi-scale image features from the FPN, including $F_{\text{Cam.}}\in\mathbb{R}^{(B_{S}\times N_{C})\times C_{\text{Cam}}\times\frac{H}{8}\times\frac{W}{8}}$ for richer semantic information and another at a reduced resolution of $\frac{H}{16},\frac{W}{16}$. We opt to utilize the feature with the resolution $\frac{H}{8},\frac{W}{8}$ due to its more comprehensive semantic content.
@@ 3.2 GlobalAlign
Figure 4: An overview of GlobalAlign pipline. The GlobalAlign module addresses the issue of misalignment in LiDAR-camera BEV feature fusion. During training, we add offset noise to simulate the global misalignment problem in the camera and LiDAR BEV features. It is supervised through a simple CBR-module to learn the offsets of camera BEV features. During testing, we do not introduce noise and employ learnable offsets for forward inference.
@@
In the real world, feature misalignment due to calibration matrix discrepancies between LiDAR and camera sensors is inevitable. While the LocalAlign module mitigates local misalignment issues in the camera-to-BEV process, deviations may still exist in camera BEV features. During LiDAR-camera BEV fusion, despite being in the same spatial domain, inaccuracies in depth from the view transformer and overlooking of global offsets between LiDAR BEV and camera BEV features lead to global misalignment.
@@
To tackle the global misalignment issue described above, we introduce the GlobalAlign module, employing learnable offsets to achieve alignment of global multi-modal BEV features. As shown in Figures 2 and 4, we use clean dataset such as nuScenes [4] for training, which exhibit minimal deviations that can be considered negligible. The supervised information is derived from the features obtained after the fusion and convolution of LiDAR and camera BEV features. During training, we introduce global offset noise and employ learnable offsets. In the LiDAR branch, the LiDAR feature is flattened along the Z-axis to form the LiDAR BEV feature, defined as $F^{L}_{B}\in\mathbb{R}^{B_{S}\times\hat{C_{L}}\times H_{B}\times W_{B}}$. Initially, we concatenate $F^{L}_{B}$ and $F^{C}_{B}$ to obtain a fused BEV feature, denoted as $F^{MM}_{B}\in\mathbb{R}^{B\times(\hat{C_{C}}+\hat{C_{L}})\times H_{B}\times W_{B}}$. Subsequently, $F^{MM}_{B}$ undergoes a convolution operation, resulting in a new fused feature, denoted as $\hat{F_{B}}\in\mathbb{R}^{B_{S}\times\hat{C_{L}}\times H_{B}\times W_{B}}$. It is noteworthy that $\hat{F_{B}}$ will be utilized as a supervision signal during the training process.
@@
As shown in Figure 4, we introduce random offset noise to the camera dimension of $F^{MM}_{B}$ to obtain a new noisy feature $F^{MM}_{N}\in\mathbb{R}^{B_{S}\times(\hat{C_{C}}+\hat{C_{L}})\times H_{B}\times W_{B}}$, simulating global misalignment issues originating from camera BEV features. Notably, the LiDAR BEV feature is directly flattened, thus more accurate. Then, $F^{MM}_{N}$ is input into the MM-Align module for global offset learning. $F^{MM}_{N}$ is processed through the CBR-module with basic convolution operations to learn offsets, defined as $F^{O}\in\mathbb{R}^{B_{S}\times 2\times H_{B}\times W_{B}}$, where 2 refers to the offset coordinates $(u,v)$. Subsequently, LiDAR BEV features $F^{L}_{B}$ and $F^{O}$ undergo grid sampling to generate new deform weights, defined as $F^{D}_{W}\in\mathbb{R}^{B_{S}\times\hat{C_{L}}\times H_{B}\times W_{B}}$. The purpose of grid sampling is to utilize offsets for spatial transformation of LiDAR BEV feature $F^{L}_{B}$, with learnable shifts dynamically adjusting to capture spatial dependencies more flexibly than standard convolution operations. Afterwards, $F^{D}_{W}$ is multiplied by LiDAR BEV features $F^{L}_{B}$ to dynamically adjust features, followed by standard convolution operations through the CBR-module, culminating in the output Deform BEV defined as $F^{D}_{B}\in\mathbb{R}^{B_{S}\times\hat{C_{L}}\times H_{B}\times W_{B}}$. Finally, during training, we supervise $F^{D}_{B}$ using $\hat{F_{B}}$ previously mentioned, and employ the $L_{\text{Align}}$ for supervision as follows:

### v3

# https://github.com/adept-thu/GraphBEV

@@ 1 Introduction
Figure 1: (a) Feature misalignment primarily arises from overlooking the projection matrix errors between LiDAR and camera, leading to LiDAR-to-camera providing inaccurate depth for surrounding neighbors. (b) We propose GraphBEV, enhancing LiDAR-to-camera projected depth with neighboring depths through graph-based neighbor information construction for enriched contextual depth feature learning. Subsequently, we achieve the alignment of global multi-modal features by simulating LiDAR and camera BEV features’ offsets and employing learnable offsets. (c) Empirical results reveal that our GraphBEV surpasses the BEVFusion [34] on the nuScenes by a margin of 1.6% mAP on nuScenes validation dataset [4] and by over 8.3% on noisy misalignment settings [12].
@@
3D object detection is a critical component of autonomous driving systems, designed to accurately identify and localize objects such as cars, pedestrians, and other elements within a 3D environment [50, 59]. Current practices predominantly adhere to a multi-modal fusion paradigm such as BEVFusion [34, 29] for robust and high-quality detection. Different modalities often provide complementary information. For instance, images contain rich semantic representations but lack depth information. In contrast, point clouds offer geometric and depth information but are sparse and devoid of semantic information. Therefore, effectively leveraging the advantages of multi-modal data while mitigating their limitations is crucial for enhancing the robustness of perception systems [59].
@@
Feature misalignment is a significant challenge in practical applications of multi-modal 3D object detection, primarily caused by calibration matrix errors between LiDAR and camera sensors [71, 12, 50], as shown in Figure 1(a). Multi-modal 3D object detection has evolved from early point-level [47, 56, 17, 33, 57, 69] and feature-level [53, 10, 2, 9, 51, 48] methods to the currently prevalent methods of Bird’s Eye View (BEV) fusion like BEVFusion [29, 34]. Although effective on clean datasets like nuScenes [4], the performance of BEVFusion[34] deteriorates on misaligned data, as illustrated in Figure 1(c). This performance decline is primarily due to calibration errors between LiDAR and the camera, exacerbated by factors like road vibrations [71]. These inherent errors cannot be corrected through online calibration, presenting a significant challenge[71, 12].
@@
Most feature-level multi-modal methods [24, 10, 1, 67, 74] employ Cross Attention to query features of a specific modality, circumventing the need for projection matrices. A few feature-level multi-modal methods [70, 51, 48, 9, 23, 72, 22, 19] have sought to mitigate these errors through the use of projection offsets or neighboring projections. A few BEV-based methods, such as ObjectFusion [5], eliminate the camera-to-BEV transformation during fusion to align object-centric features across different modalities. MetaBEV [14] utilizes Cross Deformable Attention for feature misalignment but overlooks depth estimation errors in view transformation, aligning features only during LiDAR and camera BEV fusion.
@@
BEVFusion[34, 29] unites camera and LiDAR data in BEV space, enhancing detection but overlooks feature misalignment in real-world applications. It is primarily evident in two aspects: 1) BEVFusion[34] transforms multi-image features into a unified BEV representation using BEVDepth’s[26] explicit depth supervision from LiDAR-to-camera. While this LiDAR-to-camera strategy offers more reliable depth than LSS[40], it overlooks the misalignment between LiDAR and camera in real-world scenarios, leading to local misalignment. 2) In the LiDAR-camera BEV fusion, misalignment of BEV features due to depth inaccuracies is overlooked by directly concatenating representations and applying basic convolution, as described in BEVFusion[34], leading to global misalignment.
@@
In this work, we propose a robust fusion framework, named GraphBEV to address the above feature misalignment problems. To tackle the local misalignment issue in camera-to-BEV transformation, we propose a LocalAlign module that first obtains neighbor depth information through a Graph in the View Transformation step of BEVFusion’s [34] camera Branch, based on the explicit depth provided by the LiDAR-to-camera projection. Subsequently, we propose a GlobalAlign module that encodes the projected depth from LiDAR-to-camera and the neighboring depth through dual depth encoding to generate a new Reliable depth representation incorporating neighbor information. Furthermore, to address the global misalignment problem in the fusion of LiDAR and camera BEV features, we effectively resolve the global misalignment between the two BEV feature representations by dynamically generating offsets. To validate the effectiveness of GraphBEV, we evaluated it on the nuScenes dataset, a well-known benchmark for 3D object detection. Our GraphBEV not only achieves state-of-the-art (SOTA) performance under clean settings but also outperforms BEVFusion by over 8.3% on non-aligned noisy settings nuScenes dataset provided by Dong et al. [12]. Notably, our GraphBEV experiences only a slight performance drop under noisy settings compared to clean settings. Notably, our GraphBEV bolsters the real-world applicability of BEV-based methods [34, 29] by addressing sensor calibration errors. Our contributions are summarized below:
@@
    By deeply analyzing the fundamental causes of feature misalignment, we propose LocalAlign and GlobalAlign modules within our GraphBEV framework to address local misalignments from imprecise depth and global misalignments between LiDAR and camera BEV features.
@@
    Extensive experiments validate the effectiveness of our GraphBEV, demonstrating competitive performance on the nuScenes dataset. Notably, GraphBEV maintains comparable performance across both clean settings and misaligned noisy conditions.
@@ 2.3 Multi-modal 3D Object Detection
Multi-modal 3D object detection refers to using data features from different sensors and integrating these features to achieve complementarity, thus enabling the detection of 3D objects. Previous multi-model methods can be coarsely classified into three types by fusion, i.e., point-level, feature-level, and BEV-based methods. Point-level [47, 56, 17, 33, 57, 69] and feature-level [53, 10, 2, 9, 51, 48, 73] typically leverage image features to augment LiDAR points or 3D object proposals. BEV-based methods [34, 29, 14, 5, 49] efficiently unify the representations of LiDAR and camera into BEV space. Although BEVFusion[34, 29] achieves high performance, they are typically tested on clean datasets like nuScenes[4], overlooking real-world complexities, especially feature misalignment, which hampers their application.
@@ 3 Method
Figure 2: The overview of GraphBEV framework. The LiDAR branch largely follows the baselines  [34, 1] to generate LiDAR BEV features. In the camera branch, we first extract camera BEV features using the proposed LocalAlign module, which aims to address local misalignment due to sensor calibration errors. Subsequently, we simulate the offset noise of LiDAR and camera BEV features, followed by aligning global multi-modal features through learnable offsets. Notably, we only add offset noise to the GlobalAlign module during training to simulate global misalignment issues. Finally, we employ a dense detection head [1] to accomplish the 3D detection task.
To address the feature misalignment issue in previous BEV-based methods [34, 29], we propose a robust fusion framework named GraphBEV. We provide an overview of our framework in Figure 2. Taking input from different sensors, including LiDAR and camera, we first apply modality-specific encoders, Swin-Transformer [32] as camera encoder and Second[62] as LiDAR encoder, to extract their respective features. Then, the camera features are transformed into the camera BEV feature through our proposed LocalAlign module, aiming to mitigate the local misalignment caused by the projection error between LiDAR and the camera in the camera-to-BEV process of previous BEV-based methods [34, 29]. The LiDAR features are then compressed along the Z-axis to represent 3D features as 2D LiDAR BEV features. Next, we propose a GlobalAlign module that further alleviates global misalignment between different modalities, including LiDAR and camera BEV features. Finally, we append detection head [1] to accomplish the 3D detection task. Our baseline is the BEVFusion[34], within which we detail the introduction of the LocalAlign and GlobalAlign modules below.
@@ 3.1 LocalAlign
Figure 3: The overview of LocalAlign pipeline. The LocalAlign module addresses Local Misalignment from LiDAR-to-camera by enhancing the camera-to-BEV Transform with neighboring depth features using a KD-Tree algorithm for nearest-neighbor relations.
To facilitate the transformation of camera features into BEV features, BEVFusion [34, 29] adopts the LSS-based[40] methods like BEVDepth [26] leverages LiDAR-to-camera to provide projected depth, thereby enabling the fusion of depth and image features. In the process of camera-to-BEV, BEVFusion [34] and BEVDepth [26] operate under the assumption that the depth information provided by LiDAR-to-camera projection is accurate and reliable. However, they overlook the complexities inherent in real-world scenarios, where most of the projection matrices between LiDAR and cameras are calibrated manually. Such calibration inevitably introduces projection errors, leading to depth misalignment—where the depths of surrounding neighbors are projected as the pixel’s depth. This depth misalignment results in inaccuracies within the depth features, causing misalignment during multi-view transformation into BEV representations. Given that LSS-based methods [40, 26] rely on depth estimation from pixel-level features with inaccuracies in detailing, this leads to local misalignment within camera BEV features. It underscores the challenges of ensuring precise depth estimation within the BEVFusion [34] and highlights the importance of robust methods to address projection errors.
@@
Input:
The indices of the projected pixels $M_{\text{Coords}}\in\mathbb{R}^{N_{P}\times 2}$.
Hyper-parameter: Number of neighbors $K_{\text{graph}}=8$.
@@ 3       Function KD-Tree ($M_{\text{Coords}}$, ${M_{\text{Coords}}}_{i}$, $K_{\text{graph}}$)
5            Compute the Euclidean distance between ${M_{\text{Coords}}}_{i}$ and $M_{\text{Coords}}$
6            Indices = argsort(distances)
7            return $M_{\text{Coords}}[1:K_{\text{graph}}]$
8
9
@@ 6            Indices = argsort(distances)
11             Neighbors = KD-Tree($M_{\text{Coords}}$, ${M_{\text{Coords}}}_{i}$, $K_{\text{graph}}$)
12            ${M_{\text{Coords}}}_{i}$ = Neighbors
@@
We employ the KD-Tree algorithm to obtain the indices of the projected pixels’ neighbors, defined as $M_{K_{\text{Coords}}}\in\mathbb{R}^{N_{P}\times K_{\text{graph}}\times 2}$, where $K_{\text{graph}}$ denotes the number of neighbors for each projected pixel. The process is outlined in Algorithm (1). It is worth noting that we simplified the process of the KD-Tree algorithm, and the code can be referred to in scipy [^1] https://github.com/minrk/scipy-1/blob/master/scipy/spatial/ckdtree.c. Then, we obtain the surrounding neighbor depth $D_{K}\in\mathbb{R}^{B_{S}\times N_{C}\times K_{\text{graph}}\times H\times W}$ by indexing $D_{S}$ with $M_{\text{Coords}}$. Then, $D_{S}$ and $D_{K}$ simultaneously enter the Dual Transform module for deep feature encoding. The shapes of $D_{S}$ and $D_{K}$ are respectively modified to $[B_{S}\times N_{C},1,H,W]$ and $[B_{S}\times N_{C},K_{\text{graph}},H,W]$ before being fed into the Dual Transform module. This module comprises straightforward components, including convolutional layers, Batch Normalization, and ReLU activations, as illustrated in Figure 3. The outcome of this process is the Dual Depth feature, denoted as $D_{\text{SK}}$, and its shape is $[B_{S}\times N_{C},C_{\text{SK}},\frac{H}{8},\frac{W}{8}]$. The camera Encoder outputs multi-scale image features from the FPN, including $F_{\text{Cam.}}\in\mathbb{R}^{(B_{S}\times N_{C})\times C_{\text{Cam}}\times\frac{H}{8}\times\frac{W}{8}}$ for richer semantic information and another at a reduced resolution of $\frac{H}{16},\frac{W}{16}$. We opt to utilize the feature with the resolution $\frac{H}{8},\frac{W}{8}$ due to its more comprehensive semantic content.
@@ 3.2 GlobalAlign
Figure 4: The overview of GlobalAlign pipeline. The GlobalAlign module addresses the issue of misalignment in LiDAR-camera BEV feature fusion. During training, we add offset noise to simulate the global misalignment problem in the camera and LiDAR BEV features. It is supervised through a simple CBR-module to learn the offsets of camera BEV features. We do not introduce noise during testing and employ learnable offsets for forward inference.
@@
In the real world, feature misalignment due to calibration matrix discrepancies between LiDAR and camera sensors is inevitable. While the LocalAlign module mitigates local misalignment issues in the camera-to-BEV process, deviations may still exist in camera-BEV features. During LiDAR-camera BEV fusion, despite being in the same spatial domain, inaccuracies in depth from the view transformer and overlooking of global offsets between LiDAR BEV and camera BEV features lead to global misalignment.
@@
To tackle the global misalignment issue described above, we introduce the GlobalAlign module, employing learnable offsets to achieve alignment of global multi-modal BEV features. As shown in Figures 2 and 4, we use clean datasets such as nuScenes [4] for training, which exhibit minimal deviations that can be considered negligible. The supervised information is derived from the features obtained after the fusion and convolution of LiDAR and camera BEV features. During training, we introduce global offset noise and employ learnable offsets. In the LiDAR branch, the LiDAR feature is flattened along the Z-axis to form the LiDAR BEV feature, defined as $F^{L}_{B}\in\mathbb{R}^{B_{S}\times\hat{C_{L}}\times H_{B}\times W_{B}}$. Initially, we concatenate $F^{L}_{B}$ and $F^{C}_{B}$ to obtain a fused BEV feature, denoted as $F^{MM}_{B}\in\mathbb{R}^{B\times(\hat{C_{C}}+\hat{C_{L}})\times H_{B}\times W_{B}}$. Subsequently, $F^{MM}_{B}$ undergoes a convolution operation, resulting in a new fused feature, denoted as $\hat{F_{B}}\in\mathbb{R}^{B_{S}\times\hat{C_{L}}\times H_{B}\times W_{B}}$. Notably, $\hat{F_{B}}$ will be utilized as a supervision signal during the training process.
@@
As shown in Figure 4, we introduce random offset noise to the camera dimension of $F^{MM}_{B}$ to obtain a new noisy feature $F^{MM}_{N}\in\mathbb{R}^{B_{S}\times(\hat{C_{C}}+\hat{C_{L}})\times H_{B}\times W_{B}}$, simulating global misalignment issues originating from camera BEV features. Notably, the LiDAR BEV feature is directly flattened, thus more accurate. Then, $F^{MM}_{N}$ is input into the MM-Align module for global offset learning. $F^{MM}_{N}$ is processed through the CBR-module with basic convolution operations to learn offsets, defined as $F^{O}\in\mathbb{R}^{B_{S}\times 2\times H_{B}\times W_{B}}$, where two refers to the offset coordinates $(u,v)$. Subsequently, LiDAR BEV features $F^{L}_{B}$ and $F^{O}$ undergo grid sampling to generate new deform weights, defined as $F^{D}_{W}\in\mathbb{R}^{B_{S}\times\hat{C_{L}}\times H_{B}\times W_{B}}$. The purpose of grid sampling is to utilize offsets for spatial transformation of LiDAR BEV feature $F^{L}_{B}$, with learnable shifts dynamically adjusting to capture spatial dependencies more flexibly than standard convolution operations. Afterward, $F^{D}_{W}$ is multiplied by LiDAR BEV features $F^{L}_{B}$ to dynamically adjust features, followed by standard convolution operations through the CBR-module, culminating in the output Deform BEV defined as $F^{D}_{B}\in\mathbb{R}^{B_{S}\times\hat{C_{L}}\times H_{B}\times W_{B}}$. Finally, during training, we supervise $F^{D}_{B}$ using $\hat{F_{B}}$ previously mentioned, and employ the $L_{\text{Align}}$ for supervision as follows: